# What’s in an embedding? Would a rose by any embedding smell as sweet?

### v1

@@ III Knowledge representation in LLMs
Recent articles from researchers in Anthropic AI and Open AI shed some light on the internal representations of commercial-grade LLMs [Roose2024NYT, templeton2024scaling, 3]. The Anthropic AI team examined one of their LLMs, Claude 3 Sonnet, using a method called ”dictionary learning” to discover patterns in the activation of neuron combinations when Claude was asked to discuss specific topics. Approximately 10 million of these patterns or features were identified.
@@
Figure 5: ”Geometric” representation of features in Claude 3 Sonnet [templeton2024scaling]
@@
In this context, some of the concepts and techniques that were developed in the era of symbolic AI in the 1970s-1980s, such as rule-based representations, frames, and ontologies that were used in expert systems, could be helpful in developing such ”analytical” approximations of real-world knowledge. Recent work shows how this could be done in different applications to develop mechanistic ”algebraic” approximations of the world to complement the ”geometric” view [4, 5, 6, 7, lenat2023LLM-CYC].
@@ IV From LLMs to LKMs for Science and Engineering
The main challenges that plague current LLMs have been around in AI since the era of expert systems. As heuristic rule-based expert systems emerged in the early 1980s - they were the ”LLMs” of that era - critiques observed their main failings such as their shallow understanding, limited generalization, brittleness of behavior, lack of reasoning transparency, poor inference, unreliability, etc. [4]. We find the same old issues resurfacing and some new ones due to the probabilistic nature of the new models [lenat2023LLM-CYC] in the context of LLMs.
@@
Although the connectionist camp seems to have won the latest battle for ideas, we believe that, in the long run, AI will be determined by the successful integration of both the symbolic and connectionist ideas. We need the symbolic paradigm to capture first-principles-based mechanistic ”deep” knowledge and reasoning, and connectionist representations for data-driven empirical knowledge [4, 5, 6, 7, lenat2023LLM-CYC]. .

### v2

@@ III Knowledge representation in LLMs
Recent articles from researchers in Anthropic AI and Open AI shed some light on the internal representations of commercial-grade LLMs [3, 4, 5]. The Anthropic AI team examined one of their LLMs, Claude 3 Sonnet, using a method called ”dictionary learning” to discover patterns in the activation of neuron combinations when Claude was asked to discuss specific topics. Approximately 10 million of these patterns or features were identified.
@@
Figure 5: ”Geometric” representation of features in Claude 3 Sonnet [4]
@@
In this context, some of the concepts and techniques that were developed in the era of symbolic AI in the 1970s-1980s, such as rule-based representations, frames, and ontologies that were used in expert systems, could be helpful in developing such ”analytical” approximations of real-world knowledge. Recent work shows how this could be done in different applications to develop mechanistic ”algebraic” approximations of the world to complement the ”geometric” view [6, 7, 8, 9, 10].
@@ IV From LLMs to LKMs for Science and Engineering
The main challenges that plague current LLMs have been around in AI since the era of expert systems. As heuristic rule-based expert systems emerged in the early 1980s - they were the ”LLMs” of that era - critiques observed their main failings such as their shallow understanding, limited generalization, brittleness of behavior, lack of reasoning transparency, poor inference, unreliability, etc. [6]. We find the same old issues resurfacing and some new ones due to the probabilistic nature of the new models [10] in the context of LLMs.
@@
Although the connectionist camp seems to have won the latest battle for ideas, we believe that, in the long run, AI will be determined by the successful integration of both the symbolic and connectionist ideas. We need the symbolic paradigm to capture first-principles-based mechanistic ”deep” knowledge and reasoning, and connectionist representations for data-driven empirical knowledge [6, 7, 8, 9, 10]. .